# A deep backward regression-based scheme for high-dimensional nonlinear partial differential equations

### 16 Mar 2026 v1

@@ 3.3 Convergence analysis
Suppose the assumptions (i)-(iv) hold. Furthermore, assume that $X_{0}\in L^{2+\delta}\left(\mathcal{F}_{0},\mathbb{R}^{d}\right)$, for some $\delta>0$, and $g$ is a $L_{g}$-Lipschitz function. Then, there exists a bounded sequence $\mathcal{K}_{i}$ (uniformly in $i,N$) such that for $\mathcal{NN}_{y,i}=\mathcal{G}_{\mathcal{K}_{i},1+d,n}^{\zeta_{\kappa},L,\mathbf{n}}$ and $\mathcal{NN}_{z,i}=\mathcal{G}^{\zeta_{\kappa},L,\mathbf{n}}_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},1+d,n\times d}$, we have
$\max\limits_{0\leq i\leq N}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\mathcal{Z}_{i}|^{2}ds\right]\leq Ch,$
with a grouping size $\kappa=O\left(2\sqrt{d}N^{3}\right)$, depth $L+1=O\left(d^{2}\right)$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=$ $O\left(\left(2\sqrt{d}N^{3}\right)^{d^{2}-1}\right)$ in the case $d>1$; if $d=1$, take $\kappa=O\left(N^{3}\right)$, depth $L+1=3$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=O\left(N^{3}\right)$; $K=O(N^{3})$, $K=\max\limits_{0\leq i\leq N}\{K_{i}\}$.
Proof. Step 1. Define
$\left\{\begin{aligned} \tilde{\mathcal{U}}^{m}_{t_{i}}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y+1,i})+hf\left(t_{i},\tilde{X}_{i}^{\pi,m},\tilde{\mathcal{U}}^{m}_{t_{i}},\tilde{\mathcal{V}}^{m}_{t_{i}}\right)\right],\\ \tilde{\mathcal{V}}^{m}_{t_{i}}&:=\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y+1,i})\frac{(\Delta W_{i}^{m})^{\top}}{h}\right],\end{aligned}\right$. (3.42)
where $\tilde{p}_{i}^{m}=(t_{i},\tilde{X}_{i}^{\pi,m})$. From [31, 33, 32], we have that there exist measurable deterministic functions $y_{i}:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ and $z_{i}:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n\times d}$, such that $\tilde{\mathcal{U}}^{m}_{t_{i}}=y_{i}(t_{i},\tilde{X}_{i}^{\pi,m})$ and $\tilde{\mathcal{V}}^{m}_{t_{i}}=z_{i}(t_{i},\tilde{X}_{i}^{\pi,m})$ almost surely. From (3.10) and (3.42), we have
$h(\mathcal{V}^{m}_{t_{i}}-\tilde{\mathcal{V}}^{m}_{t_{i}})=$ $\mathbb{E}_{i}\bigg[\Big(\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})$
$-\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\Big)(\Delta W_{i}^{m})^{\top}\bigg],$
because the Brownian increment $\Delta W_{i}^{m}$ is conditionally centered. From Cauchy-Schwarz inequality, we have the estimate
$h|\mathcal{V}^{m}_{t_{i}}-\tilde{\mathcal{V}}^{m}_{t_{i}}|^{2}\leq$ $d\mathbb{E}_{i}\bigg[\Big(\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\Big)^{2}\bigg]$ (3.43)
$-d\bigg(\mathbb{E}_{i}\Big[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\Big]\bigg)^{2}.$
From Young inequality, we have
$|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}$ $\leq(1+\gamma h)\left|\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}$ (3.44)
$\quad+h^{2}(1+\frac{1}{\gamma h})\bigg|\mathbb{E}_{i}\big[f\left(t_{i},X_{i}^{\pi,m},\mathcal{U}^{m}_{t_{i}},\mathcal{V}^{m}_{t_{i}}\right)-f\left(t_{i},\tilde{X}_{i}^{\pi,m},\tilde{\mathcal{U}}^{m}_{t_{i}},\tilde{\mathcal{V}}^{m}_{t_{i}}\right)\Big]\bigg|^{2}$
$\leq(1+\gamma h)\left|\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}$
$+3L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}_{i}\Big[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}+|\mathcal{V}^{m}_{t_{i}}-\tilde{\mathcal{V}}^{m}_{t_{i}}|^{2}+|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}\Big].$
Plugging (3.43) into (3.44), we derive
$|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}$ $\leq(1+\gamma h-3dL^{2}(h+\frac{1}{\gamma}))\left|\mathbb{E}_{i}\left[\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})\right]\right|^{2}$
$\quad+3dL^{2}(h+\frac{1}{\gamma})\mathbb{E}_{i}\left[|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})|^{2}\right]$
$\quad+3L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}_{i}\left[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}+|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}\right]$
$\leq(1+\gamma h)\mathbb{E}_{i}\left[|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})|^{2}\right]$
$\quad+3L^{2}h^{2}(1+\frac{1}{\gamma h})\mathbb{E}_{i}\left[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}+|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}\right],$
and thus for $h$ small enough:
$|\mathcal{U}^{m}_{t_{i}}-\tilde{\mathcal{U}}^{m}_{t_{i}}|^{2}$ $\leq(1+Ch)\mathbb{E}_{i}\left[|\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y,i+1})-\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i+1},\theta^{*}_{y,i+1})|^{2}\right]$
$\quad+Ch\mathbb{E}_{i}\left[|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|^{2}\right],$ (3.45)
From Lemma 3.2 in [17], we get, for i=N1i=N-1
$|y(t_{N-1},X_{N-1}^{\pi,m})-y(t_{N-1},\tilde{X}_{N-1}^{\pi,m})|^{2}$ $=|\mathcal{U}^{m}_{t_{N-1}}-\tilde{\mathcal{U}}^{m}_{t_{N-1}}|^{2}$ (3.46)
$\leq(1+Ch)L_{g}^{2}|X_{N-1}^{\pi,m}-\tilde{X}_{N-1}^{\pi,m}|^{2}.$
(3.46) shows that $y(t_{N-1},X_{N-1}^{\pi,m})$ is $(1+Ch)L_{g}^{2}$-Lipschitz with respect to $x$. Similarly, if $\mathcal{NN}_{y,i+1,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i+1},\theta^{*}_{y+1,i})$ is Lipschitz with respect to $x$, then $y(t_{i},X_{i}^{\pi,m})$ is also Lipschitz with respect to $x$.
Step 2. We approximate $y(t_{i},X_{i}^{\pi,m})$ by means of a $\mathcal{K}_{i}$-Lipschitz GroupSort neural network $\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i})$ in $\mathcal{NN}_{y,i}=\mathcal{G}_{\mathcal{K}_{i},1+d,n}^{\zeta_{\kappa},L,\mathbf{n}}$. From Lemma 3.2, we have the uniform error $2\mathcal{K}_{i}R\epsilon$ on $R,R]^{d}$. Then by the reverse discrete Gronwall inequality, one has
$|y(t_{i},X_{i}^{\pi,m})-y(t_{i},\tilde{X}_{i}^{\pi,m})|$ $\leq\mathcal{K}|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|,$
@@
$|\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{y,i})-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(\tilde{p}^{m}_{i},\theta^{*}_{y,i})|$ $\leq\mathcal{K}|X_{i}^{\pi,m}-\tilde{X}_{i}^{\pi,m}|,$
uniformly with respect to $i$ and $N$ for some constants $\mathcal{K}$. In the same way, we approximate $z(t_{i},X_{i}^{\pi,m})$ via a $\sqrt{\frac{d}{h}}\mathcal{K}_{i}$-Lipschitz GroupSort neural network $\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p^{m}_{i},\theta^{*}_{z,i})$ in $\mathcal{NN}_{z,i}=\mathcal{G}_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},1+d,n\times d}^{\zeta_{\kappa},L,\mathbf{n}}$. By Lemma 3.2, we have also the uniform error $2\frac{d}{\sqrt{h}}\mathcal{K}_{i}R\epsilon$ on $R,R]^{d}$. Therefore, $\sqrt{h}z(t,x)$ and $\sqrt{h}\mathcal{NN}_{z,t,1+d,n\times d}^{\varrho,L,\mathbf{n}}(t,x,\theta^{*}_{z})$ are $d\mathcal{K}$-Lipschitz with respect to $x$ and uniformly with respect to $N$.
Step 3. From the Hölder inequality, the Markov inequality and Lemma 3.2, we have the estimates
$\inf_{\mathcal{NN}_{y,i}\in\mathcal{G}_{\mathcal{K}_{i},1+d,n}}\mathbb{E}\left[\left|y(p^{m}_{i})-\mathcal{NN}_{y,i}(p^{m}_{i},\theta_{y,i})\right|^{2}\right]$
[…]
for $q>0$ and $2q=2+\delta$ withe $\delta$ as in the statement of the proposition. From the Lipschitz continuity of $y(t,x)$ and $\mathcal{NN}_{y,t,1+d,n}^{\varrho,L,\mathbf{n}}(p^{m}_{t},\theta^{*}_{y,t})$ with respect to $x$, we derive
[…]
From [30], we have the standard estimate
[…]
Hence, we deduce by (3.3)
$\inf_{\mathcal{NN}_{y,i}\in\mathcal{G}_{\mathcal{K}_{i},1+d,n}}\mathbb{E}\left[\left|y(p^{m}_{i})-\mathcal{NN}_{y,i}(p^{m}_{i},\theta_{y,i})\right|^{2}\right]\leq C\left(R^{2}\varepsilon^{2}+\frac{1+R^{2}\varepsilon^{2}}{R^{2}}\right).$ (3.49)
In the same way, repeating (3.3) and (3.48) by replacing $\mathcal{NN}_{y,i,1+d,n}(p_{i},\theta_{y,i}^{*})$ by $\mathcal{NN}_{z,i,1+d,n\times d}(p_{i},\theta^{*}_{z,i})$ and $y(t_{i},X_{i}^{\pi})$ by $z(t_{i},X_{i}^{\pi})$ and recalling that $\sqrt{h}z(t_{i},X_{i}^{\pi})$ and $\sqrt{h}\mathcal{NN}_{z,i,1+d,n\times d}(p_{i},\theta^{*}_{z,i})$ are $\mathcal{K}d$-Lipschitz with respect to $x$ and uniformly with respect to $N$, we have
@@
$\inf_{\mathcal{NN}_{z,i}\in\mathcal{G}_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},1+d,n\times d}}\mathbb{E}\left[h\left|z(p^{m}_{i})-\mathcal{NN}_{z,i}(p^{m}_{i},\theta_{z,i})\right|^{2}\right]\leq C\left(R^{2}\varepsilon^{2}+\frac{1+R^{2}\varepsilon^{2}}{R^{2}}\right).$ (3.50)
(3.49) and (3.50) show that to achieve a convergence rate of $O(\frac{1}{N})$ in Theorem 3.1, it suffices to choose the parameters $R,\varepsilon,K$ such that
$N^{2}R^{2}\varepsilon^{2}=O(\frac{1}{N}),\quad N^{2}\frac{1+R^{2}\varepsilon^{2}}{R^{2}}=O(\frac{1}{N}),\quad\frac{N^{2}}{K}=O(\frac{1}{N}),$ (3.51)
which is verified with $R=O(N^{\frac{3}{2}}),\varepsilon=O(N^{-3}),K=O(N^{3})$. From Lemma 3.2, we have that if $d>1$, we can choose the GroupSort neural network with a grouping size $\kappa=O\left(\lceil 2\sqrt{d}N^{3}\rceil\right)$, depth $L+1=O\left(d^{2}\right)$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=$ $O\left(\left(2\sqrt{d}N^{3}\right)^{d^{2}-1}\right)$. If $d=1$, take $\kappa=O\left(N^{3}\right)$, depth $L+1=3$, and width $\sum_{\ell=0}^{L-1}n_{\ell}=O\left(N^{3}\right)$. From $\Phi$ satisfies the Lipschitz condition, (3.14), we obtain (see [19, 36])
$\mathbb{E}\left[\left|\Phi(X_{T})-\Phi(X_{N}^{\pi,m})\right|^{2}\right]+\mathbb{E}\left[\sum\limits_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}|Z_{s}-\bar{Z}_{t_{i}}|^{2}ds\right]\leq Ch.$ (3.52)
From (3.49), (3.50), (3.51) and (3.52), one yields

### 21 May 2026 v3

@@ 3.3 Convergence analysis
Suppose that the assumptions (i)–(iv) hold. Furthermore, assume that $X_{0}\in L^{2+\delta}(\mathcal{F}_{0},\mathbb{R}^{d})$, for some $\delta>0$, and that $g$ is an $L_{g}$-Lipschitz function. Then, there exists a bounded sequence $\{\mathcal{K}_{i}\}_{i=0}^{N}$, uniformly in $i$ and $N$, such that the spatial slices of the GroupSort neural networks satisfy
$x\longmapsto\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}((t_{i},x),\theta_{y,i}^{*})\in G_{\mathcal{K}_{i},d,n}^{\zeta_{\kappa},L,\mathbf{n}},$
and
$x\longmapsto\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}((t_{i},x),\theta_{z,i}^{*})\in G_{\sqrt{\frac{d}{h}}\mathcal{K}_{i},d,n\times d}^{\zeta_{\kappa},L,\mathbf{n}}.$
Let
$K_{\mathrm{MC}}:=\min_{0\leq i\leq N-1}K_{i}$
denote the minimal number of inner Monte Carlo samples used to approximate the conditional expectations. If
$K_{\mathrm{MC}}\asymp N^{3},$
then the DBR approximation $(\mathcal{Y}_{i},\mathcal{Z}_{i})_{0\leq i\leq N-1}$ satisfies
$\max_{0\leq i\leq N}\mathbb{E}\left[\left|Y_{t_{i}}-\mathcal{Y}_{i}\right|^{2}\right]+\mathbb{E}\left[\sum_{i=0}^{N-1}\int_{t_{i}}^{t_{i+1}}\left|Z_{s}-\mathcal{Z}_{i}\right|^{2}ds\right]\leq Ch,$
where $C$ is independent of $N$, $h$, and $K_{\mathrm{MC}}$.
More precisely, the above rate is obtained by choosing the truncation radius $R$, the GroupSort approximation parameter $\eta$, and the inner Monte Carlo sample size $K_{\mathrm{MC}}$ as
$R=O\left(N^{\frac{3}{\delta}}\right),\qquad\eta=O\left(N^{-\frac{3}{2}-\frac{3}{\delta}}\right),\qquad K_{\mathrm{MC}}\asymp N^{3}.$
Consequently, in the case $d>1$, the GroupSort networks can be chosen with grouping size
$\kappa=O\left(\left\lceil 2\sqrt{d}\,N^{\frac{3}{2}+\frac{3}{\delta}}\right\rceil\right),$
depth
$L+1=O(d^{2}),$
and width
$\sum_{\ell=0}^{L-1}n_{\ell}=O\left(\left(2\sqrt{d}\,N^{\frac{3}{2}+\frac{3}{\delta}}\right)^{d^{2}-1}\right).$
If $d=1$, one can take
$\kappa=O\left(N^{\frac{3}{2}+\frac{3}{\delta}}\right),\qquad L+1=3,\qquad\sum_{\ell=0}^{L-1}n_{\ell}=O\left(N^{\frac{3}{2}+\frac{3}{\delta}}\right).$
Proof.
Step 1. We first prove the Lipschitz propagation of the target functions which will be approximated by GroupSort neural networks. For notational simplicity, for every fixed time level $t_{i}$, we write
$p_{i}^{x}:=(t_{i},x),\qquad x\in\mathbb{R}^{d}.$
Set
$\Psi_{N}(x):=g(x),$
and, for $0\leq i\leq N-1$,
$\Psi_{i}(x):=\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{x},\theta_{y,i}^{*}).$
For $x\in\mathbb{R}^{d}$, define the one-step Euler process starting from $x$ at time $t_{i}$ by
$X_{i+1}^{i,x}=x+\mu(t_{i},x)h+\sigma(t_{i},x)\Delta W_{i}.$ (62)
Then the deterministic functions $y_{i}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ and $z_{i}:\mathbb{R}^{d}\to\mathbb{R}^{n\times d}$ are defined recursively by
$y_{i}(x)$ $=\mathbb{E}\left[\Psi_{i+1}\left(X_{i+1}^{i,x}\right)\right]+hf\left(t_{i},x,y_{i}(x),z_{i}(x)\right),$ (63)
$z_{i}(x)$ $=\mathbb{E}\left[\Psi_{i+1}\left(X_{i+1}^{i,x}\right)\frac{\Delta W_{i}^{\top}}{h}\right],$
where the expectation is taken with respect to the Brownian increment $\Delta W_{i}$. In particular, by the Markov property of the Euler scheme,
$\mathcal{U}_{i}^{m}=y_{i}(X_{i}^{\pi,m}),\qquad\mathcal{V}_{i}^{m}=z_{i}(X_{i}^{\pi,m}).$
Let $x,x^{\prime}\in\mathbb{R}^{d}$, and let $X_{i+1}^{i,x}$ and $X_{i+1}^{i,x^{\prime}}$ be driven by the same Brownian increment $\Delta W_{i}$. From the Lipschitz continuity of $\mu$ and $\sigma$, we have
$\mathbb{E}\left[\left|X_{i+1}^{i,x}-X_{i+1}^{i,x^{\prime}}\right|^{2}\right]\leq(1+Ch)|x-x^{\prime}|^{2}.$ (64)
Assume that $\Psi_{i+1}$ is $\mathcal{K}_{i+1}$-Lipschitz with respect to the spatial variable. Define
$\Delta\Psi_{i+1}:=\Psi_{i+1}\left(X_{i+1}^{i,x}\right)-\Psi_{i+1}\left(X_{i+1}^{i,x^{\prime}}\right),\qquad\overline{\Delta\Psi}_{i+1}:=\mathbb{E}\left[\Delta\Psi_{i+1}\right].$
Then, by (64),
$\mathbb{E}\left[\left|\Delta\Psi_{i+1}\right|^{2}\right]\leq\mathcal{K}_{i+1}^{2}\mathbb{E}\left[\left|X_{i+1}^{i,x}-X_{i+1}^{i,x^{\prime}}\right|^{2}\right]\leq(1+Ch)\mathcal{K}_{i+1}^{2}|x-x^{\prime}|^{2}.$ (65)
We next estimate the Lipschitz constant of $z_{i}$. Since $\mathbb{E}[\Delta W_{i}]=0$, we have
$h\left(z_{i}(x)-z_{i}(x^{\prime})\right)=\mathbb{E}\left[\left(\Delta\Psi_{i+1}-\overline{\Delta\Psi}_{i+1}\right)\Delta W_{i}^{\top}\right].$ (66)
Therefore, by Cauchy-Schwarz inequality,
$h\left|z_{i}(x)-z_{i}(x^{\prime})\right|^{2}$ $\leq d\,\mathbb{E}\left[\left|\Delta\Psi_{i+1}-\overline{\Delta\Psi}_{i+1}\right|^{2}\right]$ (67)
$\leq d\,\mathbb{E}\left[\left|\Delta\Psi_{i+1}\right|^{2}\right]$
$\leq d(1+Ch)\mathcal{K}_{i+1}^{2}|x-x^{\prime}|^{2}.$
Now we estimate the Lipschitz constant of $y_{i}$. From (63),
$y_{i}(x)-y_{i}(x^{\prime})=\overline{\Delta\Psi}_{i+1}+h\Delta f_{i},$ (68)
where
$\Delta f_{i}:=f(t_{i},x,y_{i}(x),z_{i}(x))-f(t_{i},x^{\prime},y_{i}(x^{\prime}),z_{i}(x^{\prime})).$
By Young inequality,
$|a+b|^{2}\leq(1+\gamma h)|a|^{2}+\left(1+\frac{1}{\gamma h}\right)|b|^{2},$
and by the Lipschitz continuity of $f$, we obtain
$|y_{i}(x)-y_{i}(x^{\prime})|^{2}$ $\leq(1+\gamma h)\left|\overline{\Delta\Psi}_{i+1}\right|^{2}$ (69)
$\quad+3L^{2}h^{2}\left(1+\frac{1}{\gamma h}\right)\left(|x-x^{\prime}|^{2}+|y_{i}(x)-y_{i}(x^{\prime})|^{2}+|z_{i}(x)-z_{i}(x^{\prime})|^{2}\right).$
Using (66)–(67), we further get
$|y_{i}(x)-y_{i}(x^{\prime})|^{2}$ $\leq\left(1+\gamma h-3dL^{2}\left(h+\frac{1}{\gamma}\right)\right)\left|\overline{\Delta\Psi}_{i+1}\right|^{2}$ (70)
$\quad+3dL^{2}\left(h+\frac{1}{\gamma}\right)\mathbb{E}\left[\left|\Delta\Psi_{i+1}\right|^{2}\right]$
$\quad+3L^{2}h\left(h+\frac{1}{\gamma}\right)\left(|x-x^{\prime}|^{2}+|y_{i}(x)-y_{i}(x^{\prime})|^{2}\right).$
Since
$\left|\overline{\Delta\Psi}_{i+1}\right|^{2}\leq\mathbb{E}\left[\left|\Delta\Psi_{i+1}\right|^{2}\right],$
it follows from (70), for $h$ small enough, that
$|y_{i}(x)-y_{i}(x^{\prime})|^{2}$ $\leq(1+Ch)\mathbb{E}\left[\left|\Delta\Psi_{i+1}\right|^{2}\right]+Ch|x-x^{\prime}|^{2}$ (71)
$\leq\left((1+Ch)\mathcal{K}_{i+1}^{2}+Ch\right)|x-x^{\prime}|^{2}.$
Set $\mathcal{K}_{N}=L_{g}$ and define recursively
$\mathcal{K}_{i}^{2}=(1+Ch)\mathcal{K}_{i+1}^{2}+Ch,\qquad i=N-1,N-2,\ldots,0.$ (72)
By the discrete Gronwall inequality, the sequence $\{\mathcal{K}_{i}\}_{i=0}^{N}$ is bounded uniformly in $i$ and $N$. Hence,
$|y_{i}(x)-y_{i}(x^{\prime})|\leq\mathcal{K}_{i}|x-x^{\prime}|.$ (73)
Moreover, since $\mathcal{K}_{i}^{2}\geq(1+Ch)\mathcal{K}_{i+1}^{2}$, from (67) we also have
$\sqrt{h}\,|z_{i}(x)-z_{i}(x^{\prime})|\leq\sqrt{d}\,\mathcal{K}_{i}|x-x^{\prime}|.$ (74)
Thus, $y_{i}$ is $\mathcal{K}_{i}$-Lipschitz and $z_{i}$ is $\sqrt{d/h}\mathcal{K}_{i}$-Lipschitz with respect to the spatial variable, uniformly in $i$ and $N$.
Step 2. We approximate the spatial function $y_{i}$ by means of a $\mathcal{K}_{i}$-Lipschitz GroupSort neural network
$\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{x},\theta_{y,i}^{*})\in G_{\mathcal{K}_{i},1+d,n}^{\zeta_{\kappa},L,\mathbf{n}}.$
Since $t_{i}$ is fixed at each time step, Lemma 2 is applied to the spatial variable $x\in\mathbb{R}^{d}$. Hence, on $R,R]^{d}$, there exists such a network satisfying
$\sup_{x\in[-R,R]^{d}}\left|y_{i}(x)-\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{x},\theta_{y,i}^{*})\right|\leq 2\mathcal{K}_{i}R\varepsilon.$
Similarly, by (74), $z_{i}$ can be approximated by a $\sqrt{d/h}\mathcal{K}_{i}$-Lipschitz GroupSort neural network
@@
$\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{x},\theta_{z,i}^{*})\in G_{\sqrt{d/h}\mathcal{K}_{i},1+d,n\times d}^{\zeta_{\kappa},L,\mathbf{n}},$
and we have
@@
$\sup_{x\in[-R,R]^{d}}\sqrt{h}\left|z_{i}(x)-\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{x},\theta_{z,i}^{*})\right|\leq 2\sqrt{d}\,\mathcal{K}_{i}R\varepsilon.$
Step 3. We now estimate the approximation errors on the whole space. Let
$q:=1+\frac{\delta}{2},\qquad 2q=2+\delta.$
For notational simplicity, write
$\mathcal{N}_{y,i}(x;\theta_{y,i}):=\mathcal{NN}_{y,i,1+d,n}^{\varrho,L,\mathbf{n}}(p_{i}^{x},\theta_{y,i}),\qquad\mathcal{N}_{z,i}(x;\theta_{z,i}):=\mathcal{NN}_{z,i,1+d,n\times d}^{\varrho,L,\mathbf{n}}(p_{i}^{x},\theta_{z,i}),$
where $p_{i}^{x}=(t_{i},x)$. Let $\mathcal{N}_{y,i}^{*}$ and $\mathcal{N}_{z,i}^{*}$ be the GroupSort networks constructed in Step 2. Since $t_{i}$ is fixed at each time step, the GroupSort approximation result is applied to the spatial variable $x\in\mathbb{R}^{d}$. Hence, for $R\geq 1$, there exists a GroupSort network $\mathcal{N}_{y,i}^{*}$ such that
$\sup_{x\in[-R,R]^{d}}\left|y_{i}(x)-\mathcal{N}_{y,i}^{*}(x)\right|\leq CR\eta,$ (75)
where $\eta>0$ is the approximation parameter in Lemma 2. Similarly, there exists a GroupSort network $\mathcal{N}_{z,i}^{*}$ such that
$\sup_{x\in[-R,R]^{d}}\sqrt{h}\left|z_{i}(x)-\mathcal{N}_{z,i}^{*}(x)\right|\leq CR\eta.$ (76)
Here and below, the constant $C$ may depend on $d,n,T,L$, and on the uniform bound of the sequence $\{\mathcal{K}_{i}\}_{i=0}^{N}$, but is independent of $N,R,\eta$, and the Monte Carlo sample sizes.
We first estimate the approximation error of the $Y$-component. By splitting the domain into $\{|X_{i}^{\pi,m}|\leq R\}$ and $\{|X_{i}^{\pi,m}|>R\}$, we have